OpenStudy (anonymous):

What is the value of : \(72 \times 12 \div 6(5 + 1)\)

OpenStudy (turingtest):

is this some kind of order of operations trick or something?

OpenStudy (anonymous):

I am in doubt I want to clear that...

OpenStudy (anonymous):

You will use PEMDAS here ??

OpenStudy (turingtest):

yeah the parentheses first either way\[72\times12\div6\times6\]now the OO doesn't matter, but I'm not sure whether to consider it as\[72\times12\div(6\times6)\]

OpenStudy (amistre64):

its just mathed from left to right since ther eis no trump on a / or *

OpenStudy (anonymous):

You should do multiply first there @TuringTest

OpenStudy (turingtest):

yeah, I'm inclined to agree with that^ just\[72\times12\div6\times6\]

OpenStudy (turingtest):

it won't matter, you can do the parentheses last if you want

OpenStudy (anonymous):

Yes what will you use here division or multiplication in priority ??

OpenStudy (anonymous):

According to PEMDAS you should multiply first 72 with 12 right?? @TuringTest

OpenStudy (anonymous):

u go with BODMAS rule 72*12/6*6 72*2*6 864

OpenStudy (amistre64):

pemdas doesnt have a division multiplication priority ....

OpenStudy (anonymous):

@Nick_Black I will come to BODMAS later..

OpenStudy (anonymous):

So what will you do next @amistre64

OpenStudy (amistre64):

i already stated that i would follow the protocol and work it left to right

OpenStudy (amistre64):

1/3x is spose to be 1/3*x but laziness has us forgo the ()

OpenStudy (anonymous):

So according to you @amistre64 864 \(\div\) 6 \(\times 6\) = 864 right ??

OpenStudy (amistre64):

72*12 = a a/6 = b b*(6) = finale

OpenStudy (anonymous):

Yes I did the same..

OpenStudy (amistre64):

then yes :)

OpenStudy (amistre64):

without a proper use of grouping symbols; thats the way i would read it -- cold turkey

OpenStudy (anonymous):

I am in doubt because : BODMAS gives priority to Division.. And PEMDAS gives to Division.. Still there is no difference in answers why how ??

OpenStudy (anonymous):

Sorry PEMDAS to multiplication...

OpenStudy (amistre64):

pedmas doesnt give any priority

OpenStudy (anonymous):

We have to just go left to right ??

OpenStudy (amistre64):

yes, in the absense of any groupings, we go left to right

OpenStudy (anonymous):

Yeah we have to solve brackets first...

OpenStudy (amistre64):

have to? no but its helpful

OpenStudy (anonymous):

For our Convenience...

OpenStudy (amistre64):

i like to format it as: 1) groupings 2) dm or md 3) +- or -+ as prioritites

OpenStudy (anonymous):

Okay I understand that..

OpenStudy (anonymous):

More suggestion will be appreciated..

OpenStudy (anonymous):

Thank you @Nick_Black Your medal is remaining.. I will remember it..

OpenStudy (amistre64):

we should take its derivative while we have it on the table :)

OpenStudy (anonymous):

Can you @amistre64 give him a medal??

OpenStudy (anonymous):

What ??? @amistre64

OpenStudy (amistre64):

i can, aaaannnndddd my medals are worth more to the smartscore .... if that matters

OpenStudy (anonymous):

Yes I can understand your point in that..

OpenStudy (anonymous):

thanks to u i knew about PEMDAS today

OpenStudy (anonymous):

That is why I posted it...

OpenStudy (anonymous):

Because I am also from India..

OpenStudy (anonymous):

@amistre64 I have one more doubt if you can clear that..

OpenStudy (amistre64):

i can try, but i tend to give more doubts than i clear up lol

OpenStudy (anonymous):

What is FOIL..?? I have heard everyone saying that.

OpenStudy (amistre64):

FOIL is a memory aid to those who have lost the ability to multiply

OpenStudy (anonymous):

Is it having any full form ??

OpenStudy (anonymous):

In case of distribution it is used I guess..

OpenStudy (amistre64):

take 2 binomial expressions (a+b) and (c+d) to multiply them we simply a+b c+d ------ ac + bc ad + bd -------------- ac + bc + ad + bd FOIL is used to express the exact same thing, but thru theu use of memory aids to get you to multiply the parts togther in a specific order: Firsts, Outers, Innards, Lasts

OpenStudy (amistre64):

since the memory aid is only useful to binomials; i feel its rather useless overall ...

OpenStudy (anonymous):

Okay thank you so much, I understood now @amistre64

OpenStudy (amistre64):

youre welcome; the aggravation occurs when people ask then; how do i foil these? towit the memory aid becomes useless clutter :/

OpenStudy (anonymous):

Ha ha ha...